OpenStudy (bakonloverk):

a triangle has an area of 75 square centimeters. if the base and height have equal length what is this length? round to the nearest tenth if necessary.

OpenStudy (anonymous):

The equation for the area of a triangle is \[A = 1/2 bh\] So set up that equation with the values you are given such that \[75=1/2 (2b)\] Since both base and height are equal you can simplify b*h to 2b or 2h and then just solve for the variable.

OpenStudy (anonymous):

Area=(1/2)*base*height base=length, so Area= (1/2)*length*length = 75 cm^2 solve for length length*length=2*75 cm^2 so length=sqrt(150 cm^2)

OpenStudy (bakonloverk):

im still confused

OpenStudy (anonymous):

no, it isn't 2b, it is b^2 or l^2

OpenStudy (anonymous):

^what he said, careless mistake :p sorry

OpenStudy (bakonloverk):

how do i set up the problem?

OpenStudy (anonymous):

Area=(1/2)*base*height base=length, so Area= (1/2)*length*length = 75 cm^2 solve for length length*length=2*75 cm^2 so length=sqrt(150 cm^2) length = 150^(1/2) cm

OpenStudy (anonymous):

Use the information you are given and plug it into the equation for a triangle\[A = \frac{ 1 }{ 2 } base * height\]and since the base and height are equal you can simplify it to \[75 = \frac{ 1 }{ 2 } * x^2\] and solve for x.

OpenStudy (bakonloverk):

never mind i got the answer